HUBBLE SPACE TELESCOPE DAILY REPORT #968 PERIOD COVERED: 0800 EST Daily Status Report as of 252/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 3584 (Probing the Voids) The High Resolution Spectrograph was used to make observations of the active galactic nuclei of Markarian galaxy IZWI which is located behind an extra-galactic void. This is to study the nature, abundance, and distribution of Hydrogen I absorbing clouds. The on-board target acquisition was successful. Most of the observations were reported to look okay. 1.2 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.3 Completed WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) The Wide Field Camera was used to make Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. The images were received and reported to look okay. 1.4 Completed FOC/96 4153 (FOC UV Throughput Monitoring) The Faint Object Camera was used to make observations of the ultraviolet standard star BPM16274 to help determine the ultraviolet throughput of the HST-FOC. The observations were executed as scheduled, and no problems were reported . 1.5 Completed FOS/BL 5053 (Cycle 2 FOS Location of Spectra: Y-Base Maps) The Faint Object Spectrograph (Blue Side) was used to make a Y-Base calibration test, spectra of the internal Pt-Ne lamps looking through the 0.3 arc-second apertures for every grating. These observations mapped the face of the photocathod using 24 Y-steps and 1 X-step. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 6 Successful : 6 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.2 Complete FOC/96 3801 (Searching for QSOs Suitable for Subsequent Observations of HE II) 3.3 Complete FOS/RD 3676 (Quasar Absorption Line Studies of Starburst Galaxy Environment) 3.4 Complete Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.5 Complete PC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) 3.6 Complete Two Sets of PC 5073 (WF/PC Cycle 3 Calibration: PC Flats Wide. 2) 3.7 Complete Two Sets of PC 4381 (HST Medium - Deep Survey - Cycle 3) 3.8 Complete Two Sets of PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) 3.9 Complete FOS/BL 5046 (Cycle 3 FOS Photometric Monitoring) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-